Evaluate the integral of cos(sqrt(x))dx.


1
Expert's answer
2021-08-03T16:01:48-0400

Let us evaluate the integral "\int\cos(\sqrt x)dx".

Let us use the transformation "t=\sqrt x." Then "x=t^2," and hence "dx=2tdt."

It follows that

"\int\cos(\sqrt x)dx=2\int t\cos tdt"


"|u=t,dv=\cos tdt,\ du=dt, v=\sin t|"


"=2t\sin t-2\int \sin t dt=2t\sin t+2\cos t +C"


"=2\sqrt x\sin (\sqrt x)+2\cos (\sqrt x) +C."